# Efficiently Ranking Software Variants with Minimal Benchmarks

### v1

@@ 4.1. Bisection Sampling
Obviously, we would like to sample more cleverly than pure random. For that, we need to improve on two points: reducing the sampling variance and eliminating tests faster than one at a time.
@@
The idea is to draw on the following fact: If we add more tests to a solution to the RTSM problem, then it is also a solution. Then, conversely, if we have a set of tests that is not a solution, none of the subsets is a solution. Given a solution $T$ for each test $t\in T$ we can check if $solves_{T}(T\setminus\{t\})$ is truetrue, that is, if with all the information except for test $t$ we have a solution. If $solves_{T}(T\setminus\{t\})$ is falsefalse then it must be that the information of test $t$ is necessary if we want to reduce $T$ further; therefore, it is pointless to try to remove it. Iterating this process over all tests of $T$ partitions the tests into two sets: a set of tests $N$ that are necessary, and a set of tests $P$ that can potentially be removed. This process is described in Algorithm 1. Notice that it is quite costly because we need to do this for every test that can be potentially removed, but in practice, we hope that it reduces the number of elements we have to sample from.
@@
1:for $t\in P$ do
2:  $C\leftarrow(N\cup P)\setminus\{t\}$
3:  if not $solves_{\mathcal{T}}(C)$ then
4:    $N\leftarrow N\cup\{t\}$
5:    $P\leftarrow P\setminus\{t\}$
6:  end if
7:end for
8:return $N,P$
@@
The idea is to use a bisection algorithm as in (Bentley et al., 2019), for example, where a bisection algorithm is used to find the relevant code locations. Bisection algorithms are similar to binary search; the general idea is to use a divide-and-conquer approach where the search space is split, then tested, and based on the results of the test, it is reduced, and then the process repeats until one element is found. Here we are looking for one set among a set of test sets. Here we have $P$, the set of tests from which we want to sample, $N$, the set of tests that are necessary, and $\mathcal{T}$ the test suite considered, but the set of possible solutions is too large, making the bisection algorithm impractical. Instead, the idea is to randomly divide $P$ into two sets $A$ and $B$ of the same size. Then we can do the following: check if $A$ or $B$ is a solution with $solves_{\mathcal{T}}(A\cup N)$, if that is the case, we divide the size of $P$ by a factor of two, and we can repeat the process. If none of them are solutions, then we make the following assumption that either $A$ or $B$ is necessary; therefore, we compute the solution where $A$ is necessary and the solution where $B$ is necessary and return the best. In each step, we divide the size of $P$ by two, which allows much faster progress.
@@
This process is formalized in Algorithm 2. The first step is to use Algorithm 1 to reduce the variance, and then we can proceed with the sampling.
@@
2:if $P$ is empty then
3:  return NN
@@
7:if $solves_{\mathcal{T}}(A\cup N)$ then
8:  return $BS(A,N,\mathcal{T})$
9:else if $solves_{\mathcal{T}}(B\cup N)$ then
10:  return $BS(B,N,\mathcal{T})$
11:else
12:  $S_{A}\leftarrow BS(B,A\cup N,\mathcal{T})$
13:  $S_{B}\leftarrow BS(A,B\cup N,\mathcal{T})$
14:  return best of $S_{A}$ or $S_{B}$
15:end if
@@ 4.2. Divide and Conquer
In reality, $solves_{K\cup P}(K)$ does a linear regression, which complexity is O($|\mathcal{A}||K|^{2}|P|)$ where $\mathcal{A}$ is the set of variants. Therefore, testing for a large test set $K$ becomes very costly, as in $FindNecessary$, where we remove a single test. Thus, a more scalable approach needs to be used.
@@
The idea is to use a divide-and-conquer approach and to use bisection sampling as a subroutine. Fig. 2 illustrates at a high level our divide-and-conquer approach with our iterative strategy. In the first step, we split $\mathcal{T}$ into multiple smaller sets and solve these smaller instances of the RTSM problem. From then on, at each iteration, we will merge two solutions of the previous step. We do this simply by considering the instance of the RTSM problems, where we keep only the tests in either solution. This is a new instance of the RTSM problem that we solve. We keep iterating until there is only one instance left. In the worst case, in all of the sub-instances, we could not find better solutions than including every test of the sub-instance; therefore, the last step is similar to not doing any divide and conquer, but in all other cases, we potentially solve smaller problems, enabling faster solving. The initial split thus has a large variance and, therefore, can have a large impact on the performance of this approach.
The process is described in detail in Algorithm 3. But before diving into the explanations of the algorithm, let us convince ourselves that this process gives a valid solution. Given two solutions $A$ and $B$ for the RTSM instances $T_{A}$ and $T_{B}$, is $S$ a solution to the RTSM instance $A\cup B$ also a solution to the entire RTSM instance $T_{A}\cup T_{B}$? The answer is no, the reason being that when we build the predictions, we build them on the exact true data, so when we consider only $A\cup B$ we may be able to remove the tests because we can predict them accurately up to their ranks but not accurately enough for use in the prediction of other tests. To do this correctly at minimal cost, when a solution $S$ is found to the RTSM instance $A\cup B$ we check if it is also a solution to the entire RTSM instance $T_{A}\cup T_{B}$, if it is we accept it and continue otherwise we try again a certain number of times until we arbitrarily set $S$ to $A\cup B$.
Algorithm 3 DC($\mathcal{T}$, $n$, SampleSample): from a set of tests $\mathcal{T}$ by splitting it into $n$ parts and using the SampleSample to sample solutions, find a solution by using divide and conquer
@@
5:  $S\leftarrow Sample(A\cup B,\emptyset,A\cup B)$
6:  if not $solves_{T_{A}\cup T_{B}}(S)$ then $\triangleright$ check if $S$ is NOT a valid merge
@@ 5. Other Approaches
We describe a few methods inspired by related work and intuition that we consider as baselines. They can also be used as divide-and-conquer subroutines, including iterative solving, algorithms, and will try solutions and test them with solvessolves.
@@
Random search is a simple baseline that can be easily adapted to the RTSM problem. The random search involves randomly sampling a set of tests and then checking if the set forms a solution that can be returned at any time. Rather than sampling among all possible candidates, the process can be adapted to be iterative. Initially, there are $n$ tests in the suite, then we randomly search for a set of tests of size n1n-1.
@@
Another simple black-box baseline in a greedy search approach. At each step, we can try to remove the test that provides the least information; since all tests have the same weight, we chose to remove the test with the lowest metric variance because, on average, it should contribute the least to the ranking. Note that we also tried with the lowest mean metric, and the results were approximately the same, so we decided not to report this variant for brevity reasons.
@@
The WRTSM problem can be transformed into an MILP problem, except that since our decision variables are the presence or absence of a test in the solution, this disables the ability to also optimize $W$, which weights the tests in the subsets. Let us use the notations introduced in Definition 3.3: $\mathcal{A}$ is the set of variants, $\mathcal{T}$ is the set of tests, $p_{k,i}$ if the performance of variant $k$ in the test $i$, $cost(t_{i})$ is the cost of the test $t_{i}$ and the weights of the solution $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$$\in\mathbb{R}^{+}$ for the test $t_{i}$. Let us explain our encoding; the presence or absence of the test $t_{i}\in\mathcal{T}$ in the solution is encoded as ${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\in\{0,1\}$. Our constraints are encoded as follows:
@@
which translates to the ranking should be the same. Thus, our objective is to minimize:
@@
However, this formulation does not take into account the weights of the tests in the solution. They can be included in the following way into our constraints for variant $k$ and variant $l$:
@@
Notice that this is no longer a linear program, since both the $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} and the $\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}$} are decision variables. Therefore, we use the previous linear version even though it solves a simpler version of the problem, but with less agency; therefore, the solutions of the MILP approaches are not optimal for the problem where the $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} can be set, although they are optimal for the ${\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}=1$ cases. Observe that this formulation also only enables us to consider a Kendall of exactly 1, if we wanted to consider a Kendall of $T$, then we would use the following formulation for variant $k$ and variant $l$:
@@
since ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=1$ for a concordant pair and ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=-1$ for a discordant pair, translating exactly the left-hand side into the definition of the Kendall. This means that changing the Kendall also implies that this is not a linear program anymore, thus MILP is a very rigid approach. Note that MILP is used as a standalone baseline without divide-and-conquer. In our implementation, MILP uses the COIN (CBC) solver (John Forrest et al., 2024).
@@
Principal Component Analysis (PCA) is a technique used to reduce dimensionality while minimizing loss in terms of variance. We use PCA on the performance matrix and look at the most important vector, the one who explains most of the variance. Then we choose to remove the test that is least important for this vector, that are the one whose coefficient is the least correlated with this vector.
@@ 5.1. Failed Attempts
Based on racing algorithms such as irace (López-Ibáñez et al., 2016), we adapt their approach based on Friedman’s test (and, 1937) and produce the same decisions irace would take. In other words, it works iteratively and chooses a test to remove at each step. The test chosen is based on Friedman’s test used in irace to detect the poorest performers; here, our intuition is that it should detect the test that has the least discriminative power and remove it. We implemented it in such a way that if we run irace on our setup, except that it does not have new data at each step, then irace would remove the same tests in the same order. Applying it in practice over all our benchmarks led to no cost reduction at all on all seeds; therefore, this baseline is discarded.
@@
There is a large body of work on the minimization of the test suite that uses genetic algorithms (Harman and Jones, 2001; Yoo and Harman, 2012). However, most of these approaches are white- or gray-box approaches. Unlike these approaches, there are neither features nor associated code; therefore, we cannot adapt their approaches. We can still adapt the approach of genetic algorithms to our case. Our genes consist of zeros and ones; a one indicates that we keep a test in the test suite, and a zero indicates that it is discarded. Then, we can frame the RTSM problem as a multi-objective optimization problem by considering two independent objectives, namely 1) minimizing the size of the test suite and 2) minimizing the ranking error so that it stays below the desired threshold. Genetic algorithms, unlike other methods, cannot guarantee that the ranking error stays below the desired threshold; that is why we have to consider it as an objective function to minimize it. In order for genetic algorithms to work with our two objectives, we use the following formula:
@@ 6. Experiments
  • RQ4: How does our method compare to the others?
  • RQ5: How does the different components affect performance?

### v2

@@ 4.1. Bisection Sampling
Obviously, we would like to sample more intelligently than pure random subsets of $\mathcal{T}$. For that, we need to improve on two points: reducing the sampling variance and eliminating tests faster than one at a time.
@@
The idea is to draw on the following fact: If we add more tests to a solution to the RTSM problem, then it is also a solution. Then, conversely, if we have a set of tests that is not a solution, none of the subsets is a solution. Given a solution $T$ for each test $t\in T$ we can check if $\text{findWeightsAndCheck}_{T}(T\setminus\{t\})$ is true, that is, if with all the information except for test $t$ we have a solution. If $\text{findWeightsAndCheck}_{T}(T\setminus\{t\})$ is false then it must be that the information of test $t$ is necessary if we want to further reduce $T$; therefore, it is pointless to remove it. Iterating this process over all tests of $T$ partitions the tests into two sets: a set of tests $N$ that are necessary, and a set of tests $P$ that can potentially be removed. This process is described in Algorithm 1. Notice that it is quite costly because we need to do this for every test that can be potentially removed, but in practice, we hope that it reduces the number of elements we have to sample from. It has complexity O($|\mathcal{A}||N\cup P|^{2}|P|)$.
@@
[…]
3:for $t\in P$ do
4:  $C\leftarrow(N^{\prime}\cup P^{\prime})\setminus\{t\}$
5:  if not $\text{findWeightsAndCheck}_{\mathcal{T}}(C)$ then
6:    $N^{\prime}\leftarrow N^{\prime}\cup\{t\}$
7:    $P^{\prime}\leftarrow P^{\prime}\setminus\{t\}$
8:  end if
9:end for
10:return $N^{\prime},P^{\prime}$
@@
The idea is to use a bisection algorithm as in (bisection19), for example, where it is used to find the relevant code locations. Bisection algorithms are similar to binary search; the general idea is to use a divide-and-conquer approach where the search space is split, then tested, and based on the results of the test, it is reduced, and then the process repeats until one element is found. Here we are looking for one set among a set of test sets. Here we have $P$, the set of tests from which we want to sample, $N$, the set of tests that are necessary, and $\mathcal{T}$ the test suite considered, but the set of possible solutions is too large, making the bisection algorithm impractical. Instead, the idea is to randomly divide $P$ into two sets $A$ and $B$ of the same size. Then we can do the following: check if $A$ or $B$ is a solution with $\text{findWeightsAndCheck}_{\mathcal{T}}(A\cup N)$, if that is the case, we divide the size of $P$ by a factor of two, and we can repeat the process. If none of them are solutions, then we make the following assumption: $A$ or $B$ is necessary; therefore, we compute the solution where $A$ is necessary and the solution where $B$ is necessary and return the best. In each step, we divide the size of $P$ by two, which allows much faster progress.
@@
This process is formalized in Algorithm 2. The first step is to use Algorithm 1 to reduce the variance, and then we can proceed with the sampling. Notice that in any recursive call to $BS$, the size of $P$ is divided by 2, therefore the longest call chain we can have is $O(\log|P|)$. In the worst case scenario, we make two recursive calls, therefore the worst case complexity is $O(2^{\log|P|}|\mathcal{A}||N\cup P|^{2}|P|)=O(|\mathcal{A}||N\cup P|^{2}|P|^{2})$. It is expected that it only has polynomial complexity since we do not try every possible combination, we use the bisection to make large cuts in order to provide solutions faster.
@@
2:if $P^{\prime}$ is empty then $\triangleright$ Every test is necessary
3:  return $N^{\prime}$
@@
7:  return $BS(A,N^{\prime},\mathcal{T})$
8:else if $\text{findWeightsAndCheck}_{\mathcal{T}}(B\cup N^{\prime})$ then
9:  return $BS(B,N^{\prime},\mathcal{T})$
[…]
11:  $S_{A}\leftarrow BS(B,A\cup N^{\prime},\mathcal{T})$
12:  $S_{B}\leftarrow BS(A,B\cup N^{\prime},\mathcal{T})$
13:  return best of $S_{A}$ or $S_{B}$
14:end if
@@ 4.2. Divide and Conquer
As mentioned previously, $\text{findWeightsAndCheck}_{K\cup P}(K)$ has already quadratic complexity, and our $BS$ algorithm is in $O(|\mathcal{A}||N\cup P|^{2}|P|^{2})$ and since initially $P=\mathcal{T}$ and $N=\emptyset$ then the worst case complexity is $O(|\mathcal{A}||\mathcal{T}|^{4})$. Therefore, a more scalable, practical approach must be used.
@@
The idea is to use a divide-and-conquer approach and to use sampling, here BS, as a subroutine. Fig. 2 illustrates at a high level our divide-and-conquer approach with our iterative strategy. In the first step, we split $\mathcal{T}$ into multiple smaller sets and sample solutions for these smaller instances of the RTSM problem. From then on, at each iteration, we merge two solutions of the previous step. By merging, we mean that the RTSM instances are merged, but instead of starting from scratch we start as a base solution from the union of the two solutions found at the previous step. We keep iterating until there is only one instance left. The goal is that, when facing larger instances, the set $P$ of instances to potentially remove is much smaller reducing the complexity. In the worst case, in all the sub-instances, no better solutions than including every test of the sub-instance could be found; therefore, the last step is similar to not doing any divide and conquer. The initial split thus has a large variance and, therefore, can have a large impact on the performance of this approach.
The process is described in detail in Algorithm 3. But before diving into the explanations of the algorithm, let us convince ourselves that this process gives a valid solution. Given two solutions $A$ and $B$ for the RTSM instances $T_{A}$ and $T_{B}$, is $S$ a solution to the RTSM instance $A\cup B$ also a solution to the entire RTSM instance $T_{A}\cup T_{B}$? The answer is no, the reason being that when we use solvesolve, we perform a linear regression that depends on the exact true data. When we consider only $A\cup B$, we may be able to remove the tests, because we have the exact data to get the correct ranks, but the predicted performance cannot be fed as true data for another linear regression; it may be too inaccurate. To do this correctly at minimal cost, when a solution $S$ is found to the RTSM instance $A\cup B$ we check if it is also a solution to the entire RTSM instance $T_{A}\cup T_{B}$, if it is we accept it and continue otherwise we try again a certain number of times until we arbitrarily set $S$ to $A\cup B$.
Algorithm 3 DC($\mathcal{T}$, $n$, BSBS): from a set of tests $\mathcal{T}$ by splitting it into $n$ parts and using $BS$ a sampling strategy by default ours, find a solution by using divide and conquer
@@
5:  $S\leftarrow BS(A\cup B,\emptyset,A\cup B)$
6:  if not $\text{findWeightsAndCheck}_{T_{A}\cup T_{B}}(S)$ then $\triangleright$ check if $S$ is NOT a valid merge
@@ 5. Other Approaches
We describe a few methods inspired by related work and intuition that we consider as baselines. Notice that Algorithm 3 takes $BS$ as input, it implies that if we have sampling strategies, they can also be used in combination with this approach.
@@
Random search is a simple baseline that can be easily adapted to the RTSM problem. The random search involves randomly sampling a set of tests and then checking if the set forms a solution that can be returned at any time. Rather than sampling among all possible candidates, the process can be adapted to be iterative. Initially, there are $n$ tests in the suite, then we randomly search for a set of tests of size n1n-1, if a valid solution is found then we sample a set of tests of size n2n-2 from the solution of size n1n-1 and so on. The complexity is $O(k|\mathcal{A}|n^{2}$) where $k$ is the number of samples taken in the worst case.
@@
Another simple black-box baseline in a greedy search approach. At each step, we can try to remove the test that provides the least information; since all tests have the same weight, we chose to remove the test with the lowest metric variance because, on average, it should contribute the least to the ranking. Note that we also tried with the lowest mean metric, and the results were approximately the same, so we decided not to report this variant for brevity reasons. More formally, we remove a $t^{*}\in\text{argmin}_{t}\text{Var}(P_{*t})$ where $P_{*t}$ is the vector containing the performance of all the variants on $t$. We break ties by alphabetical order. The complexity is $O(|\mathcal{A}||\mathcal{T}|^{2}$) in the worst case.
@@
We can transform the WRTSM problem into an approximate MILP formulation. Let us use the notations introduced in Definition 3.3: $\mathcal{A}$ is the set of variants, $\mathcal{T}$ is the set of tests, $p_{k,i}$ is the performance of variant $k$ on test $i$, $cost(t_{i})$ is the cost of test $t_{i}$ and the weights of the solution $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$$\in\mathbb{R}^{+}$ for the test $t_{i}$. Let us explain our encoding; the presence or absence of the test $t_{i}\in\mathcal{T}$ in the solution is encoded as ${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}}\in\{0,1\}$. Our constraints to keep the same ranking are encoded as follows:
@@
Thus, our objective is to minimize:
@@
However, this formulation does not take into account the weights $W\in\mathbb{R}^{|\mathcal{T}|\times 1}$ of the tests in the solution. They can be included in the following way in our constraints for variant $k$ and variant $l$:
@@
Notice that this is no longer a linear program, since both $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} and $\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}u_{i}$} are decision variables. Therefore, we use the previous linear version, without the weights $W\in\mathbb{R}^{|\mathcal{T}|\times 1}$. This formulation solves a simpler version of the problem; therefore, the solutions of the MILP approaches are not optimal for the problem where the $\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}$} can be set, although they are optimal for the ${\color[rgb]{1,0.265625,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.265625,0.19921875}w_{i}}=1$ cases. Observe that this formulation also only enables us to consider a Kendall of exactly 1, if we wanted to consider a Kendall of $T$, then we would use the following formulation for variant $k$ and variant $l$:
@@
since ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=1$ for a concordant pair and ${\color[rgb]{0.2265625,0.5859375,0.0078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2265625,0.5859375,0.0078125}c_{k,l}}=-1$ for a discordant pair, translating exactly the left-hand side into the definition of the Kendall. This means that changing the target Kendall also implies that this becomes a non linear program, thus MILP is a very rigid approach and we must use the limited original formulation. Note that MILP is used as a standalone baseline without divide-and-conquer. In our implementation, MILP uses the COIN (CBC) solver (John2024coin).
@@
Principal Component Analysis (PCA) is a technique used to reduce dimensionality while minimizing loss in terms of variance. We use PCA on the performance matrix and look at the most important vector, $v\in\mathbb{R}^{\mathcal{T}}$, the one who explains most of the variance. Then we choose to remove one test that is the least important for this vector: $t^{*}\in\text{argmin}_{t}v_{t}$, in case of ties, we take the smallest test by alphabetical order.
@@ 5.1. Failed Attempts
Based on racing algorithms such as irace (LOPEZIBANEZ201643), we adapt their approach based on Friedman’s test (Friedman01121937) and produce the same decisions irace would take. In other words, it works iteratively and chooses a test to remove at each step. The test chosen is based on Friedman’s test used in irace to detect the poorest performers; here, our intuition is that it should detect the test that has the least discriminative power and remove it. We implemented it in such a way that if we run irace on our setup, except that it does not have new data at each step, then irace would remove the same tests in the same order. Applying it in practice over all our benchmarks led to no cost reduction at all on all seeds; therefore, this baseline is discarded.
@@
There is a large body of work on the minimization of the test suite that uses genetic algorithms (harman2001search; harman2012RT). However, most of these approaches are white- or gray-box approaches. Unlike these approaches, there are neither features nor associated code; therefore, we cannot adapt their approaches. We can still adapt the approach of genetic algorithms to our case. Our genes consist of zeros and ones; a one indicates that we keep a test in the test suite, and a zero indicates that it is discarded. Then, we can frame the RTSM problem as a multi-objective optimization problem by considering two independent objectives, namely 1) minimizing the size of the test suite and 2) minimizing the ranking error so that it stays below the desired threshold. Genetic algorithms, unlike other methods, cannot guarantee that the ranking error stays below the desired threshold; that is why we have to consider it as an objective function to minimize it. In order for genetic algorithms to work with our two objectives, we use the following formula:
@@ 6. Experiments
  • RQ4: How many benchmark runs are needed to break even between minimization cost and savings?
  • RQ5: How does our method BISS compare to the others?
  • RQ6: How do the different components affect performance?